for every 15 boys who atendend the program there were 10 girls. how many girls were there if there were 60 boys who atandend in program?
Answer by gonzo(654)

the ratio of boys to girls is 15/10.
15/10 is the same as 3/2.
---
if b = boys,
and g= girls,
and r = ratio,
then:
r = b/g
---
since they gave you the number of boys, and you want to find the number of girls, and you know what the ratio is, then you can use this equation to find the number of girls.
---
the equation is:
r = b/g
when r = 3/2, and b = 60, this equation becomes:
3/2 = 60 / g
if you multiply both sides of this equation by g, you get:
g * (3/2) = 60
if you multiply both sides of this equation by 2, you get:
2*g * (3/2) = 2 * 60
which reduces to:
g * 3 = 120
if you divide both sides of this equation by 3, you get:
g = 120/3 = 40
---
number of boys is 60.
number of girls is 40.
ratio of boys to girls is 60/40.
---
60/40 is the same as 6/4 is the same as 3/2 which is the ratio of boys to girls that we started with.
---
your answer is:
40 girls attended if 60 boys attended.
---